# Two-photon interference between mutually-detuned resonance fluorescence signals scattered off a semiconductor quantum dot

### v2

@@ Abstract
To conclude, we have experimentally investigated post-selective TPI between mutually detuned RF signals scattered off a single QD. By invoking the model that treats all the RF photons as result of spontaneous emission (Eq. 2), it is possible to find full correspondence between the predictions of the model and the experimental results. Beyond validating the theoretical model, these findings are also highly relevant to quantum information processing, as we demonstrated that indistinguishable photons can be generated by an emitter, such as a quantum dot, irrespective of the excitation detuning.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grants 62250710162, 12494600, 12494604 and Beijing Natural Science Foundation under grant IS23011.
Appendix A Model for a coherently-driven two-level system
Fig. 6: Alternative pictures for a two-level emitter under weak, detuned coherent driving. a, Elastic/passive scattering; b, Spontaneous emission into temporal modes that are phase-locked by the excitation laser. $\nu_{0}$: Transition gap between the ground (|$\ket{g}$) and exited (|$\ket{e}$) states; $\Delta$: Excitation detuning.
As depicted in Fig. 6a, a TLE under a weak coherent excitation is usually considered as a passive scatterer 20, as the scattered light inherits the driving laser’s energy and linewidth (coherence). On the other hand, the scattered signal is known to exhibit photon anti-bunching over the time scale of the TLE’s radiative lifetime, which can be several orders of magnitude shorter than the driving laser’s coherence time. This sharp contrast led to the term of “sub-natural” 21 linewidth, and the reconcilliation could be understood, with careful elaboration, in terms of the interference between the passively scattered photons and the actively emitted photons resulting from actual photon absorption and re-emission 22.
We have recently proposed a simple model 16 that treats all RF photons systematically as an active result of absorption and re-emission. This model successfully accounts for both the RF’s first-order coherence and photon anti-bunching under strict resonant condition. Extending to a detuned excitation, it immediately predicts that all RF photons from the same TLE remain identical to one another irrespective to the excitation detuning, even when the RF fields may exhibit vastly different central frequencies. Fig. 6b schematically summarizes the RF model under a detuned excitation: the properties of the single photons are defined by the two-level emitter but the temporal modes they radiate into are phase-locked by the excitation laser.
For strict resonant excitation, we have previously derived the pure-state description of Eq. 2 when considering the TLE and its photon emission jointly 16, although the TLE and the emission separately are mixed states. Following similar derivation steps, here we generalize the pure-state description to cover the presence of excitation detuning. Our starting point is the model 23 of a two-level system in the frame that rotates with the driving laser and in the CW regime, i.e.,
@@ Appendix A Model for a coherently-driven two-level system
where $h=1$, $\sigma=|g\rangle\langle e|$ ($\sigma^{\dagger}=|e\rangle\langle g|$) is the atom lowering (raising) operator, $\sigma_{z}=|g\rangle\langle g|-|e\rangle\langle e|$ is the Pauli $z$ matrix, $\Omega$ is the Rabi frequency and $\Delta:=\nu-\nu_{0}$ is the excitation detuning. $\phi$ is the phase of the dipole matrix element $\mathbf{p}_{ge}=|\mathbf{p}_{ge}|e^{i\phi}$. Since it does not produce an observable effect, we let $\phi=\pi/2$.
@@
with $\gamma_{\parallel}$ being the spontaneous emission rate, $\gamma^{*}$ the pure dephasing rate of the emitter, $\rho_{a}$ is the density matrix of a two-level atom, i.e., TLE and $\mathcal{D}$ the dissipation superoperator given by
@@
$\rho_{gg}(\infty)=\frac{2\gamma_{\parallel}\left(\gamma_{\perp}^{2}+4\Delta^{2}\right)+\gamma_{\perp}\Omega^{2}}{2\gamma_{\parallel}\left(\gamma_{\perp}^{2}+4\Delta^{2}\right)+2\gamma_{\perp}\Omega^{2}},$ (11)
@@
[…]
@@
With the hypothesis that the driving laser is perfectly filtered from the radiation emitted by the source, we can leverage the relation of proportionality between the source dipole and the field operator of the collected light, i.e., $a{\color[rgb]{0,0,0}\propto}\sigma$. Therefore, we might as well let $a=B\sigma$, where $B$ is a proportional coefficient. In this way, we always have that the state of light normalized by the intensity corresponds to the steady-state of the emitter.
@@
After obtaining the reduced density matrices of the TLE and fluorescence individually, we can infer the density matrix of the TLE-photon joint system. Under the equilibrium condition of continuous-wave laser driving, at any given time $t$, the density matrix state vector of the joint system of the two-level atom and fluorescence only needs to consider $|g0\rangle$, $|e0\rangle$, and $|g1\rangle$, while excluding higher-order terms such as $|e1\rangle$ and $|g2\rangle$. This is primarily based on exciting experimental observations and theoretical results, where $g^{(2)}(0)=0$ is always valid, regardless of whether the excitation is strong or weak. Therefore, the instantaneous atom-photon density matrix can be written in the form
$\rho_{ap}=\left(\begin{array}[]{lll}\rho_{g0,g0}&\rho_{g0,g1}&\rho_{g0,e0}\\ \rho_{g1,g0}&\rho_{g1,g1}&\rho_{g1,e0}\\ \rho_{e0,g0}&\rho_{e0,g1}&\rho_{e0,e0}\end{array}\right).$ (19)
In the absence of pure dephasing, i.e., $\gamma^{*}=0$, we obtain the pure-state form of Eq. 2 and
$p_{0}=\frac{{\color[rgb]{0,0,0}\gamma_{\parallel}}^{2}+4\Delta^{2}}{{\color[rgb]{0,0,0}\gamma_{\parallel}}^{2}+4\Delta^{2}+2\Omega^{2}}=\frac{1}{1+2\eta_{ab}\bar{n}},$ (20)
$p_{1}=1-p_{0}=\frac{2\eta_{ab}\bar{n}}{1+2\eta_{ab}\bar{n}},$ (21)
where $\frac{\Omega^{2}}{{\color[rgb]{0,0,0}\gamma_{\parallel}}^{2}+4\Delta^{2}}=\eta_{ab}\bar{n}$ and $\eta_{ab}$ represents the photon absorption efficiency at weak pump limit.
Appendix B Derivation of $g^{(2)}_{\parallel}(\tau)$ under steady-state condition ($\tau\gg T_{1}$)
In this section, we will discuss the coincidence counts for different incident configurations depicted in Fig. 3 from the perspective of the input photon states. We start from the pure-state description of Eq. 2. Tracing out the TLE component state (|$\ket{g}$, |$\ket{e}$), the remaining photon sub-system is in a mixed state, which is equivalent to the pure photon state |$\ket{\psi}_{ph}=\sqrt{p_{0}}\ket{0}+\sqrt{p_{1}}e^{i2\pi\nu t}\ket{1}$ attenuated by a channel of 50% loss. Conveniently, channel losses do not affect the analysis of the first-order or second-order correlation functions as we have previously proved 16, so we are able to use the pure photon state in the subsequent derivation.
@@ Appendix B Derivation of $g^{(2)}_{\parallel}(\tau)$ under steady-state condition ($\tau\gg T_{1}$)
$=\frac{1}{4}p_{1}^{2}+\frac{1}{4}p_{1}^{2}+\left|\frac{1}{2}(p_{0}p_{1}e^{-i2\pi 2\nu t}(e^{-i2\pi 2\nu_{a}\tau}-e^{-i2\pi 2\nu_{b}\tau})|0000\rangle\right|^{2}+\frac{1}{4}\times 2(p_{1}^{3}p_{0}+p_{1}^{3}p_{0}+p_{1}^{4})$
@@
where the form of photon states (e.g., |$\ket{0000}$) follows the convention that the first two digits correspond to the $t$ and $t+\tau$ time instants of port $a$ respectively, and the last two digits correspond to the $t$ and $t+\tau$ instants of port $b$. Within the final equation, the first and second terms correspond to cases (i) and (iv) illustrated in Fig. 3, while the last term accounts for the interference between the two indistinguishable paths as shown in Fig. 3 cases (ii) and (iii). Considering finite photon indistinguishability, we obtain the second-order HOM correlation function by normalizing to the coincidence probability of $p_{1}^{2}$ at $\tau=\infty$:
@@
Fig. 7: Measurement of the RF’s first-order coherence. a, Purpose-built AMZI with a differential delay of 5 ns; b, Normalized count rates at ports $c$ (black) and $d$ (red) under a single laser excitation of $\Delta$ = 0 GHz and $\bar{n}$ = 0.002. The phase drift is caused by varying the temperature of the AMZI.
Appendix C The RF’s first-order correlation function
In measuring $g^{(1)}$, we use a second, purpose-built AMZI with a differential delay of $\tau=5$ ns as shown in Fig. 7a. This delay is much greater than the QD relaxation time but much shorter than the laser coherence time. Under this condition of $T_{1}\ll\tau\ll T_{L}$, the RF’s first-order correlation function can be written as 24, 16
$|g^{(1)}(\tau)|=\sqrt{\mathcal{M}}p_{0},$ (26)
from which we can bound photon indistinguishability to $\mathcal{M}\leq|g^{(1)}|^{2}$ using inequality $p_{0}<1$. Fig. 7b shows the interference fringes, from which we extract the first-order correlation function value of $|g^{(1)}|=0.991\pm 0.008$. This result places an upper bound of 0.98 for $\mathcal{M}$.
Appendix D Modeling the 0-delay HOM results
@@
Each experimental $\mathcal{V}(0)$ value is calculated from the corresponding HOM measurements of $g^{(2)}_{\perp}(0)$ and $g^{(2)}_{\parallel}(0)$ using to Eq. 1. However, the measured $g^{(2)}_{\perp}(0)$ values at large detunings fall below the theoretical minimum of 0.5 for completely distinguishable but intensity-balanced RF inputs. To model this irregularity ($g^{(2)}_{\perp}<0.5$), we introduce an empirical parameter $x$ ($0\leq x\leq 1$) to quantify the intensity-imbalance between the RF streams. We fit the measured HOM correlation data using
@@
Among simulation parameters, $x$ is a fitting parameter and excitation detuning dependent. The auto-correlation functions $g^{(2)}(\tau)$ used are obtained from the de-convolution of the measured HBT data with the detector response $R(\tau)$. We use fixed values of $\mathcal{M}_{\perp}=0$ and $\mathcal{M}_{\parallel}=0.94$ for all excitation fluxes. $R(\tau)$ is a Gaussian function of 59 ps FWHM.

### v3

@@ Abstract
To conclude, we have experimentally investigated post-selective TPI between mutually detuned RF signals scattered off a single QD. By invoking the model that treats all the RF photons as a result of spontaneous emission (Eq. 2), it is possible to find full correspondence between the predictions of the model and the experimental results. Beyond validating the theoretical model, these findings are also highly relevant to quantum information processing, as we demonstrated that indistinguishable photons can be generated by an emitter, such as a quantum dot, irrespective of the excitation detuning.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under grants 12494600, 12494604 and 62250710162, the Beijing Natural Science Foundation under grant IS23011, and the Beijing Postdoctoral Research Foundation.
Author Contributions
#These authors contributed equally to this work.
Appendix A Theoretical modelling
A.1 Model for a coherently-driven two-level system
[…]
As depicted in Fig. 6a, a TLE under a weak coherent excitation is usually considered as a passive scatterer 27, as the scattered light inherits the driving laser’s energy and linewidth (coherence). On the other hand, the scattered signal is known to exhibit photon anti-bunching over the time scale of the TLE’s radiative lifetime, which can be several orders of magnitude shorter than the driving laser’s coherence time. This sharp contrast led to the term of “sub-natural” 28 linewidth, and the reconciliation could be understood, with careful elaboration, in terms of the interference between the passively scattered photons and the actively emitted photons resulting from actual photon absorption and re-emission 29.
We have recently proposed a pure-state model 17 in which all RF photons are systematically treated as an active result of absorption and re-emission. This model successfully accounts for both the RF’s first-order coherence and photon anti-bunching under strict resonant conditions. Extending to a detuned excitation, it immediately predicts that all RF photons from the same TLE remain identical to one another irrespective to the excitation detuning, even when the RF fields may exhibit vastly different central frequencies. Fig. 6b schematically summarizes the RF model under a detuned excitation: the properties of the single photons are defined by the two-level emitter but the temporal modes they radiate into are phase-locked by the excitation laser.
For strict resonant excitation, we have previously derived the pure-state description of Eq. 2 when considering the TLE and its photon emission jointly 17, although the TLE and the emission separately are mixed states. Following similar steps, we now generalize this pure-state description to include the case of finite excitation detuning. Our starting point is the standard model 30 of a coherently driven two-level system, expressed in the laser-rotating frame and under continuous-wave (CW) excitation:
@@ A.1 Model for a coherently-driven two-level system
where $\hbar=1$, $\sigma=|g\rangle\langle e|$ ($\sigma^{\dagger}=|e\rangle\langle g|$) is the atom lowering (raising) operator, $\sigma_{z}=|g\rangle\langle g|-|e\rangle\langle e|$ is the Pauli $z$ matrix, $\Omega$ is the Rabi frequency and $\Delta:=\nu-\nu_{0}$ is the excitation detuning with $\nu$ and $\nu_{0}$ denoting the laser and the atom transition frequencies, respectively. $\phi$ is the phase of the dipole matrix element $\mathbf{p}_{ge}=|\mathbf{p}_{ge}|e^{i\phi}$. For simplicity, and without affecting the generality of the analysis, we take $\phi=\pi/2$. All theoretical quantities ($\Omega$, $\Delta$, and $\gamma$) are expressed in angular frequency units (rad s-1), whereas the corresponding parameters in the Main Text are quoted as cyclic frequencies (Hz).
@@
with $\gamma_{\|}$ being the spontaneous emission rate, $\gamma^{*}$ the pure dephasing rate of the emitter, $\rho_{a}$ is the density matrix of a two-level atom, i.e., TLE and $\mathcal{D}$ the dissipation superoperator given by
@@
$\rho_{gg}(\infty)=zy{\frac{2\gamma_{\parallel}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+\gamma_{\perp}\Omega^{2}}{2\gamma_{\parallel}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+2\gamma_{\perp}\Omega^{2}}},$ (11)
@@
$\hat{\rho}_{a}(t\rightarrow\infty)=\frac{1}{2\gamma_{\|}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+2\gamma_{\perp}\Omega^{2}}\{$ ${\left[2\gamma_{\|}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+\gamma_{\perp}\Omega^{2}\right]|g\rangle\langle g|+\gamma_{\|}\left(\gamma_{\perp}+i\Delta\right)\Omega|g\rangle\langle e|}$ (14)
@@
With the hypothesis that the driving laser is perfectly filtered from the radiation emitted by the source, we can leverage the relation of proportionality between the source dipole and the field operator of the collected light, i.e., $a\propto\sigma$. Therefore, we might as well let $a=B\sigma$, where $B$ is a proportional coefficient. In this way, we always have that the state of light normalized by the intensity corresponds to the steady-state of the emitter.
@@
$\hat{\rho}_{p}(t\rightarrow\infty)=\frac{1}{2\gamma_{\|}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+2\gamma_{\perp}\Omega^{2}}\{$ ${\left[2\gamma_{\|}\left(\gamma_{\perp}^{2}+\Delta^{2}\right)+\gamma_{\perp}\Omega^{2}\right]|0\rangle\langle 0|+\gamma_{\|}\left(\gamma_{\perp}+i\Delta\right)\Omega|0\rangle\langle 1|}$ (18)
After obtaining the reduced density matrices of the TLE and fluorescence individually, we can infer the density matrix of the TLE-photon joint system. Under the equilibrium condition of continuous-wave laser driving, at any given time $t$, the density matrix state vector of the joint system of the two-level atom and fluorescence only needs to consider $|g0\rangle$, $|e0\rangle$, and $|g1\rangle$, while excluding higher-order terms such as $|e1\rangle$ and $|g2\rangle$. This is primarily based on exciting experimental observations and theoretical results, where $g^{(2)}(0)=0$ is always valid, regardless of whether the excitation is strong or weak. In the absence of pure dephasing, i.e., $\gamma_{\perp}=\gamma_{\|}/2$, the instantaneous atom-photon density matrix can be inferred to have the form
$\hat{\rho}_{ap}=\left(\begin{array}[]{ccc}\rho_{g0,g0}&\rho_{g0,g1}&\rho_{g0,e0}\\ \rho_{g1,g0}&\rho_{g1,g1}&\rho_{g1,e0}\\ \rho_{e0,g0}&\rho_{e0,g1}&\rho_{e0,e0}\end{array}\right)=\frac{1}{\gamma_{\|}^{2}+4\Delta^{2}+2\Omega^{2}}\left(\begin{array}[]{ccc}\gamma_{\parallel}^{2}+4\Delta^{2}&(\gamma_{\parallel}+2i\Delta)\Omega&(\gamma_{\parallel}+2i\Delta)\Omega\\ (\gamma_{\parallel}-2i\Delta)\Omega&\Omega^{2}&\Omega^{2}\\ (\gamma_{\parallel}-2i\Delta)\Omega&\Omega^{2}&\Omega^{2}\end{array}\right).$ (19)
The specific structure of $\hat{\rho}_{ap}$—with all nonzero elements phase-coherent and rank one—reveals that the joint atom–photon system remains in a pure quantum state. Consequently, the above density matrix can be equivalently expressed in the pure-state form
@@
where $\theta^{\prime}=\arctan\left(\frac{2\Delta}{\gamma_{\|}}\right)$ and
$p_{0}$ $=\frac{\gamma_{\parallel}^{2}+4\Delta^{2}}{\gamma_{\parallel}^{2}+4\Delta^{2}+2\Omega^{2}},\ p_{1}=\frac{2\Omega^{2}}{\gamma_{\parallel}^{2}+4\Delta^{2}+2\Omega^{2}}.$ (21)
By transforming back from the laser-rotating frame to the laboratory frame, we obtain the exact form of Eq. 2 for the joint TLE–photon system,
[…]
where $\theta=\arctan\left(\frac{2\Delta}{\gamma_{\parallel}}\right)+\frac{\pi}{2}$, reinstating the optical oscillation at frequency $\nu$ and the dipole phase initially set at $\pi/2$.
A.2 Coefficients $p_{0}$ and $p_{1}$
The dependencies of $p_{0}$ and $p_{1}$ on the excitation power $\bar{n}$ can be derived from the balance between photon absorption and re-emission by the TLE 17, yielding
$p_{0}=\frac{1}{1+2\eta_{ab}\bar{n}},\qquad p_{1}=\frac{2\eta_{ab}\bar{n}}{1+2\eta_{ab}\bar{n}},$ (23)
where $\eta_{ab}$ denotes the TLE’s absorption efficiency in the weak-drive regime, which depends on the excitation detuning. In the weak-drive limit, $p_{0}\!\to\!1$, indicating that the emitter predominantly remains in its ground state.
The connection between the coefficients $p_{0}$ and $p_{1}$ and the RF power spectrum can be established by evaluating the first-order coherence. Starting from the pure state in Eq. 22, the normalized first-order coherence function is
$g^{(1)}(\tau)=\frac{\langle\psi|a^{\dagger}(t)a(t+\tau)|\psi\rangle}{\langle\psi|a^{\dagger}a|\psi\rangle}.$ (24)
Evaluating this expression gives
$g^{(1)}(\tau)=p_{0}\,e^{-i2\pi\nu\tau},$ (25)
valid for $T_{1}\!\ll\!\tau\!\ll\!T_{L}$. Since the RF power spectrum is the Fourier transform of $g^{(1)}(\tau)$, the phase-coherent term $p_{0}\,e^{-i2\pi\nu\tau}$ corresponds to a narrow, laser-like spectral component centered exactly at the excitation frequency $\nu$ with a relative weight $p_{0}$, whereas the remaining fraction $p_{1}$ gives rise to the broadband background.
The laser-like and broadband fractions in the RF spectrum are conventionally referred to as the coherent and incoherent scattering components, respectively 31. Their normalized spectral weights are usually expressed as
$I_{\mathrm{coh}}=\frac{1}{1+s},\qquad I_{\mathrm{incoh}}=\frac{s}{1+s},$ (26)
where $s=\frac{2\Omega^{2}/\gamma_{\|}^{2}}{1+(2\Delta/\gamma_{\|})^{2}}$ is the saturation parameter. By identifying $s=2\eta_{ab}\bar{n}$, Eq. 26 becomes identical to Eq. 23, leading to $p_{0}=I_{\mathrm{coh}}$ and $p_{1}=I_{\mathrm{incoh}}$. This equivalence establishes a one-to-one correspondence between the conventional mixed-state RF description and the pure-state formulation we have introduced, where the joint emitter–photon system remains in a coherent superposition of the vacuum and one-photon states.
A.3 $g^{(2)}_{\parallel}(\tau)$ under steady-state condition ($\tau\gg T_{1}$)
[…]
attenuated by an effective channel with 50% loss. Conveniently, channel losses do not affect the analysis of the first-order or second-order correlation functions as we have previously proved 17, so we are able to use the pure photon state in the subsequent derivation.
@@ A.3 $g^{(2)}_{\parallel}(\tau)$ under steady-state condition ($\tau\gg T_{1}$)
[…]
@@
A.4 Modeling the 0-delay HOM results
@@
Each experimental $\mathcal{V}(0)$ value is calculated from the corresponding HOM measurements of $g^{(2)}_{\perp}(0)$ and $g^{(2)}_{\parallel}(0)$ using to Eq. 1. However, the measured $g^{(2)}_{\perp}(0)$ values at large detunings fall below the theoretical minimum of 0.5 for completely distinguishable but intensity-balanced RF inputs. To account for this irregularity ($g^{(2)}_{\perp}<0.5$), we introduce an empirical parameter $x$ ($0\leq x\leq 1$). We fit the measured HOM correlation data using
@@
Among simulation parameters, $x$ is a fitting parameter and excitation detuning dependent. The auto-correlation functions $g^{(2)}(\tau)$ used are obtained from the deconvolution of the measured HBT data with the detector response $R(\tau)$. We use fixed values of $\mathcal{M}_{\perp}=0$ and $\mathcal{M}_{\parallel}=0.94$ for all excitation fluxes. $R(\tau)$ is a Gaussian function of 59 ps FWHM.
Appendix B Further experimental details
B.1 Pulse-carving the excitation lasers
Fig. 7: Red and blue detuned excitation laser waveforms.
Two continuous-wave lasers are modulated into square-wave signals with a period of approximately 1190 ns using two independent acousto-optic modulators (AOMs), each providing an extinction ratio of about 50 dB. The AOMs are driven by synchronized trigger signals from the same function generator, with a relative phase shift of half a period between them. The chopped square pulses are detected by a photodetector and recorded on an oscilloscope, as shown in Fig. 7. It is observed that the two lasers alternate with a switching period of 595 ns. The electrical 20–80% edge times are measured as $\tau_{\rm r}\approx$24 ns (rise) and $\tau_{\rm f}\approx$30 ns (fall), respectively. These values satisfy $\tau_{\rm r},\tau_{\rm f}\gg T_{1}$ and $1/\tau_{\rm r,f}\ll\Delta_{\max}=4\,$GHz, suppressing edge-induced spectral artifacts while keeping the cross-talk between adjacent tones below $\text{1--2}\%$. In addition, the energy detuning between the two square waves is determined by measuring the beat frequency between the two lasers using a frequency counter.
B.2 Radiative lifetime $T_{1}$ under cavity resonance
The radiative lifetime $T_{1}$ is a key parameter characterizing the device performance. In this study, we use $T_{1}$ directly to calculate the excitation flux $\bar{n}$. To measure $T_{1}$, we employ a time-resolved setup using an 80 MHz picosecond mode-locked laser to resonantly excite the QD 24. The measured data are presented in Fig. 8. The time-resolved spectrum exhibits a pronounced laser background around zero delay, followed by an exponential decay component attributed to the QD emission. By fitting this decay, we extract a radiative lifetime of $T_{1}=74$ ps. In this measurement, the QD is tuned into resonance with the microcavity.
Fig. 8: Time-resolved resonance fluorescence measurement when the QD is tuned into resonance with the cavity. The black dots represent the experimental data, and the green line shows a single-exponential fit, yielding a radiative lifetime of $T_{1}=74$ ps.
B.3 Fine-structure splitting (FSS)
Fig. 9: Time-resolved PL spectrum of the QD detuned from the cavity mode. The red dashed line highlights the oscillation period.
Fig. 10: Dual-color HOM interference traces over an extended delay range under two different mutual detunings: $\bm{\Delta=0}$ GHz (left column) and $\bm{\Delta=0.5}$ GHz (right column). a, b, HOM interference traces for the parallel (blue) and orthogonal (red) polarization settings; c, d, Magnified view of the data shown in panels a and b.
The neutral exciton nature of the QD transition is confirmed by the presence of fine-structure splitting (FSS) 32. To measure the splitting, we employ cross-polarized resonant excitation, with the excitation polarization chosen such that both exciton eigenstates are equally addressed. Due to the FSS, the temporal evolution of the collected signal exhibits oscillations arising from the energy difference between the two polarsation states 33. By extracting the oscillation period, we determine the fine-structure splitting $\Delta_{\text{FSS}}$. To better resolve the oscillations, we detune the QD transition off the cavity resonance by raising the device temperature. The resulting temporal dynamics are presented in Fig. 9, where the oscillation period, highlighted by the red dashed line, is $\sim 1150$ ps, corresponding to $\Delta_{\text{FSS}}\approx 0.91$ GHz.
B.4 HOM results at the AMZI delay
In Fig. 2 of the main text, we plotted the dual-color HOM interference over a delay range of $\pm 400$ ns. To ensure completeness, Fig. 10 presents the same dataset over an extended delay range, revealing the correlation features at the characteristic delays of $\pm 595$ ns. A zoomed-in view around +595+595 ns is also included for clarity.
Here, the 595 ns delay significantly exceeds the mutual coherence time of the excitation laser sources ($T_{c}\approx 200$ ns), thereby erasing any phase correlation between the photon pairs. Consequently, no interference effect is observable at this delay. As shown in Figs. 10c and 10d, regardless of whether the excitation laser is detuned or the polarization is matched, the observed correlation converges to a dip value of 0.75. This quantitatively matches the results observed in single-photon streams under incoherent excitation conditions 23.
It is also worth noting that the delay configuration of the AMZI used here is fundamentally different from that in our prior work 17, where the AMZI delay (4.92 ns) was much shorter than the coherence time of the excitation laser. In that case, interference effects manifested prominently at the side peaks of the HOM interference traces. These rich experimental features were discussed in detail in Ref. 17.
B.5 Quantifying the attainable photon indistinguishability
Fig. 11: The RF’s first-order interference. a, Purpose-built AMZI with a differential delay of 5 ns; b, Normalized count rates at ports $c$ (black) and $d$ (red) under a single laser excitation of $\Delta$ = 0 GHz and $\bar{n}$ = 0.002. The phase drift is caused by varying the temperature of the AMZI.
In measuring $g^{(1)}$, we use a second, purpose-built AMZI with a differential delay of $\tau=5$ ns as shown in Fig. 11a. This delay is much greater than the QD relaxation time but much shorter than the laser coherence time. Under this condition of $T_{1}\ll\tau\ll T_{L}$, the RF’s first-order correlation function can be written as 34, 17
$|g^{(1)}(\tau)|=\sqrt{\mathcal{M}}p_{0},$ (33)
from which we can bound photon indistinguishability to $\mathcal{M}\leq|g^{(1)}|^{2}\times(1+2\bar{n})^{2}$ using Eq. 3 and inequality $\eta_{ab}\leq 1$.
Fig. 11b shows the interference fringes, measured with an excitation flux of $\bar{n}=0.002$ under strictly resonant excitation. We extract a first-order correlation value of $|g^{(1)}|=0.991\pm 0.008$, which sets an upper bound of $\mathcal{M}\leq 0.990$ for the attainable indistinguishability.